OpenStudy (anonymous):

what is equivalent to 5/10=1/__

jimthompson5910 (jim_thompson5910):

hint: divide 5 and 10 by the GCF of 5 and 10

OpenStudy (anonymous):

okay thanks . @jim_thompson5910

OpenStudy (anonymous):

so whats the gcf @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what factors go into 5 and 10?

OpenStudy (anonymous):

1/2

jimthompson5910 (jim_thompson5910):

yeah ultimately 5/10 = 1/2

OpenStudy (anonymous):

so what about 1/2=6/__

jimthompson5910 (jim_thompson5910):